What is the mass of the air contained in a room 6m X 5m x 4m if the pressure is 100 kPa and the temperature is 25oC


1
Expert's answer
2021-12-08T10:00:07-0500

"V = 6 \times 5 \times 4 = 120 \; m^3 = 120 \times 10^3 \;L \\

p= 100 \; kPa = 0.986923 \; atm \\

T = 25 + 273.15 = 298.15 \;K"

Ideal gas law

pV = nRT

R=0.082058 L×atm/(K×mol)

"n = \frac{pV}{RT} \\

n = \frac{0.986923 \times 120 \times 10^3 }{0.082058 \times 298.15} = 4840.70 \; mol

M(air) = 28.97 \; g/mol \\

m = 4840.70 \times 28.97 = 140235.2 \; g = 140.23 \; kg"